OpenStudy (anonymous):

Solve the system of equations using the elimination method. Enter the answer as an ordered pair. x+y=9 x-y=7

OpenStudy (callisto):

x+y=9 -(1) x-y=7 -(2) (1)+(2) (x+y) + (x-y) = 9+7 2x = 16 Solve x. Put x back to (1) and solve y. Can you do it?

OpenStudy (anonymous):

So X=4 @Callisto

OpenStudy (callisto):

Not really.. 2x = 16 Divide both sides by 2, what do you get?

OpenStudy (anonymous):

How would i solve for y though? @Callisto

OpenStudy (anonymous):

my bad x=8

OpenStudy (callisto):

So, put x=8 into (1) 8 + y = 9 Solve y. Can you do it?

OpenStudy (anonymous):

y=1 @Callisto Thankyou so much

OpenStudy (callisto):

Welcome :)